undeadknight26 (undeadknight26):

Phobee wants to show her company's profits and losses for the year. Which Microsoft program should she use? Access Excel PowerPoint Word

undeadknight26 (undeadknight26):

I think its excel!

OpenStudy (anonymous):

EXCELLLL.

OpenStudy (anonymous):

KEEP EM COMING. WOO xD

undeadknight26 (undeadknight26):

UNA MAS!!! Thomas needs to create a Microsoft Excel formula to figure out his monthly percentage of profits. Which formula should he use to figure out percentages? =12000x45% =12000/45% =12000+45% =45/12000% CREO QUE B!

OpenStudy (anonymous):

excel

OpenStudy (anonymous):

TU ES CORECTO.

undeadknight26 (undeadknight26):

HURRA!

undeadknight26 (undeadknight26):

@vlltimelow What anime is that?

OpenStudy (anonymous):

?

undeadknight26 (undeadknight26):

Your picture...what anime is it from?

OpenStudy (anonymous):

soul eater xD

undeadknight26 (undeadknight26):

Have you seen soul eater not yet? it looks interesting XD

OpenStudy (anonymous):

yes it's very interesting watch it

undeadknight26 (undeadknight26):

I still have to watch soul eater XD I started watching one piece as my first ever anime....worse choice ever...

OpenStudy (anonymous):

my first anime was soul eater xDD best choice ever :p

undeadknight26 (undeadknight26):

lol did you watch any others than the soul eaters or no?

OpenStudy (anonymous):

Oh o-o Snap.

undeadknight26 (undeadknight26):

Marr is just like wth is going on o.o

OpenStudy (anonymous):

yeah I watch blue exorcist

undeadknight26 (undeadknight26):

Oh that one was so good /.\ Rin and that sword was amazing XD

OpenStudy (anonymous):

DON'T SPOIL IT D:

OpenStudy (anonymous):

xD i dont watch anime, unless theres nothing else on, then i go to Adult swim and anime just so happens to be on xD

OpenStudy (anonymous):

anime is all i watch .-.

undeadknight26 (undeadknight26):

OMG ARE YOU NOW WATCHING IT?!

undeadknight26 (undeadknight26):

same...I've watched only like 50 or 70 ;-; yesh i has a life ;-;

OpenStudy (anonymous):

Aye, i dont judge o-o im just here, doing a history essay, NO BIG DEAL xD

OpenStudy (anonymous):

no no I'm watching spongebob at the moment :P

OpenStudy (anonymous):

OMG SPONGEBOB IS BAE. xD Thats mah SHOW <3

OpenStudy (anonymous):

XD same

OpenStudy (anonymous):

Ive seen like every episode atleast 5 times. o-o

OpenStudy (anonymous):

Soul Eater Evans is my crush :3 and Crona <3

undeadknight26 (undeadknight26):

*facepalm* Sponge bob will rot your brains -_- *sigh*

OpenStudy (anonymous):

Shush Knight D; Dont ruin my dreams.

undeadknight26 (undeadknight26):

lol american animation...

OpenStudy (anonymous):

I hate tv tbh

OpenStudy (anonymous):

especially pll it's so boring ._.

OpenStudy (anonymous):

Netflix. WOO.

OpenStudy (anonymous):

yaaaas.

undeadknight26 (undeadknight26):

Anime4psp.org animeplus.tv animeultima...thats where i live XD

OpenStudy (anonymous):

Oh, Ewww. xD

OpenStudy (anonymous):

xD guys I'll be on later and anime is never ew -.-

undeadknight26 (undeadknight26):

later! Anime anime, cars and weights for life!!!

OpenStudy (anonymous):

Oh gosh. xD